MONT 106N – Identifying Patterns Discussion Day 2 October 2, 2009

Stephanie Craig and Nick Foster

In class on October 2, 2009, you will be responsible for about 20 minutes of the class period, during which we will discuss the article “Some Funds Stop Grading on the Curve” Wall Street Journal, published on September 8, 2009. You can assume everyone will read the article before class – the information how to get it is posted on the course homepage. So you don’t need to spend a lot of time summarizing what it says. Instead, come up with several good questions that you can use to start the discussion, and then be prepared to ”moderate” (maybe answer some questions if people have not understood some point, call on people who have something to say, and so forth).

Some suggestions: Try to focus about half of the time on discussing the technical points: How are ideal histograms like the normal curve used to model possible returns from investments? How has that often been done in the past? Just what is a “fat tail distribution” and how does it differ from our standard normal curve? Then, open things up and ask: what are the real-world implications of using different assessments of investment risk? What is at stake? What should a non-expert know about stuff like this to be well-informed?

I found this by looking at a listing in Chance News 54 – see link to this on the course homepage. You can also get to the original version of the article from there. You will see discussion of the issues in Chance News that may give you a few ideas of how to get a discussion started. And of course, you can talk things over with me before class too.